Ativan 1 mg IV push the med comes in 2 mg per mL, how many mL do you give?

If cephalexin 500 mg is prescribed QID x10 days, how many grams will  the patient receive over the entire course?
Answer #1

ANSWER.

* Calculation of amount(ml) of drug required.

        Formula = Desired dose/Dose in hand * Quantity.

     Here

          Desired dose(D) = 1 mg.

         Dose in hand(H) = 2 mg.

             Quantity(Q) = 1 ml.

As per above formula

      Amount(ml) of drug required = D/H * Q

                                                      = 1/2 * 1

                                                      = 0.5 ml.

* Amount(gm) of drug patient receive the entire course.

          Ordered dose = 500 mg QID (Four times a day) * 10 days.

      So

        Total dose required over the entire course = 500 * 4 * 10

                                                                                    = 20000 mg.

                                                 1 mg = 0.001 gm.

             Total dose(gm) required = 20000 * 0.001

                                                                                   = 20 gm.